3) Prove by induction that for all natural numbers n:

(e) n2 + n is divisible by 2. (Can you prove this directly?)


1
Expert's answer
2021-05-31T14:12:45-0400

Induction on positive integers "n :"


When "n =1, n^2+n = 2" which is obviously divisible by "2."


Assume when "n = k" , that "k^2+k" is divisible by "2" .


When "n = k+1" , we have


"(k+1)^2+k+1 = k^2+2k+1+k+2"


"= (k^2+k)+2(k+1)"


From our assumption, "k^2+k" is divisible by "2" and so the whole expression is divisible by 2.


Therefore, by induction "n^2 + n" is divisible by 2 for all natural numbers n.